OpenStudy (anonymous):

Interest on $5,255 at 12% for 30 days (use ordinary interest) is

OpenStudy (anonymous):

do you remember the forumla

OpenStudy (anonymous):

i=prt?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I = (P)(r)(T)

OpenStudy (anonymous):

do you know which value is which?

OpenStudy (anonymous):

principle is 5255? rate= 12% time = 30 days ?

OpenStudy (anonymous):

correct but remember you cant use % so .12, but alta T is meaasured in years so what is 1 month out of a year

OpenStudy (anonymous):

so 1/12 ? X.12 ?

OpenStudy (anonymous):

(5255)(.12)(1/12) prt

OpenStudy (anonymous):

THANK YOU !

OpenStudy (anonymous):

what if it would have been 50 days instead of 30?

OpenStudy (anonymous):

1.6/12 i beleive since 20/30(a month) is 60%

OpenStudy (futureboy1):

630.60

OpenStudy (anonymous):

where did you get 630.60?

OpenStudy (anonymous):

? so confused lol =(

OpenStudy (futureboy1):

thats how much you will have to pay

OpenStudy (anonymous):

you cant just give answers.. what formula did you use

OpenStudy (futureboy1):

5255*0.12*1=

OpenStudy (futureboy1):

P is the principal amount, $5255.00. r is the interest rate, 12% per month, or in decimal form, 12/100=0.12. So, t is 1....month time periods. To find the simple interest, we multiply 5255 × 0.12 × 1 to get that: The interest is: $630.60

OpenStudy (anonymous):

no he needs 50 days

OpenStudy (futureboy1):

oh hold on

OpenStudy (anonymous):

1 month and .60 of a month

OpenStudy (anonymous):

over 12

OpenStudy (futureboy1):

1008.96

OpenStudy (futureboy1):

@lugerm

OpenStudy (solomonzelman):

5,255 is the entire amount/investment (or 100% of it). When you add 12% of the sum, you are really taking 12/100th of the sum. So, 12% of that would be; 5,255×12÷100 (which is =630.6). And that is daily interest once, but for 30 days, that would be 30×630.6

OpenStudy (solomonzelman):

THis is on an assumption that 12% is a daily interest, with a length of interest charge period of 30 days.